# Exploring the Boundaries of On-Device Inference: When Tiny Falls Short, Go Hierarchical

### v1

@@ 3.1. Devices
In this study, we conduct measurements on five devices shown in Fig. 2 and described in Tables 1 and 2.
@@
Arduino Nano 33 BLE Sense and ESP32 (cf. Table 1) are the least powerful devices used in this paper. On-device inference is performed using the TensorFlow Lite Micro framework, which requires loading the model as a binary file, leading to slower on-device computation. Furthermore, contrary to all other devices, it does not support Wi-fi communication. For this reason, in case of offload, the images are transmitted to the edge server via Bluetooth Low Energy (BLE).
@@
Coral Dev Board Micro (cf. Table 1) posseses higher processing power and more memory than other commodity MCUs, it can load and execute large compute-intensive DL models with higher inference accuracy. Although the device has beneficial features, in certain cases, it falls short. For example, the built-in TPU supports a limited set of instructions [^3] https://coral.ai/docs/edgetpu/models-intro/#supported-operations that directly impacts the models it can execute. The TFLite models must be compiled using edgetpu-compiler to support the built-in TPU. However, because early-exit models use an intermediate layer with an if instruction, they cannot be compiled to a model supported by the TPU because the instruction is not supported. One potential mitigation of such an issue is running the model on the CPU at the expense of increased inference latency and energy consumption.
@@
Contrary to Jetson Orin or Coral Micro, the Raspberry Pi 4B (cf. Table 2) does not benefit from any hardware accelerator. Such selection provides us with a broader view of the available system-on-chip ecosystem.
Regarding the software implementation, all the models are loaded and run using the TFLite Python API. All the necessary operations for the tested models are supported in this framework [^4] https://www.tensorflow.org/mlir/tfl_ops. This, combined with the fact that models were run on the CPU, allows for a smooth deployment even for Early Exit implementations.
@@
Jetson Orin Nano (cf. Table 2) includes an integrated GPU that features Tensor Cores. For this device, we have used TensorRT SDK [^5] https://docs.nvidia.com/deeplearning/tensorrt, which is a framework developed by NVIDIA for high-performance deep learning inference, built on top of CUDA. It optimizes the models using layer and tensor fusion and kernel tuning to maximize performance on the specific hardware of the target device. To port the pre-trained models to TensorRT, we must first convert them to ONNX. This pipeline yields correct results for base models, but for early exit we find that the TensorRT engine fails to produce a graph-capturable network in CUDA, which causes early exit models to be slower than the base implementations. To mitigate this issue, transforming the Tensorflow subgraph models to PyTorch successfully generates models that effectively utilize CUDA graphs and achieve the required performance.
Jetson Orin has two power modes: one for maximum performance (15W) with all hardware and clocks at full capacity and another for low power consumption (7W) with RAM, CPU, GPU, and other components capped. We examine both power modes to investigate their effects and observe that enabling dynamic clock rate results in higher latency and variability for relatively small energy savings. Setting the clock to maximum in each power mode results in higher performance per watt and more stable outcomes under constant load.
Figure 2. Devices used in the study. Starting from the left: Arduino Nano 33 BLE Sense, ESP32 DevKit-C, Coral Dev Board Micro, Raspberry Pi 4B, Jetson Orin Nano.
3.2. Datasets, Models and Training
We study image classification tasks using three datasets MNIST, CIFAR-10, and ImageNet-1K. As discussed in Section 1, MNIST image classification tasks can be performed on Aruduino Nano and ESP32, without offloading to a remote ML. Thus, the results for CIFAR-10 and ImageNet-1K image classification tasks are presented in the rest of the paper.
Table 1. Device characteristics of MCUs used for the measurements
@@ 3.2. Datasets, Models and Training
Table 2. Device characteristics of the moderately powerful devices used for the measurements
@@
We use TensorFlow [^6] https://www.tensorflow.org/ as the DL software library for most of this work due to its popularity and convenience in compiling models for on-device deployment, thanks to TFLite [^7] https://www.tensorflow.org/lite and TFLite Micro [^8] https://www.tensorflow.org/lite/microcontrollers. We refer to the models we introduce below as base models to differentiate them from the models we design and train using early exit branches, later in Section 5.
@@
ImageNet-1K. For ImageNet, we select ResNet-18, ResNet-50, AlexNet, and RegNetY32GF (Radosavovic et al., 2020). All these models are retrieved from PyTorch’s own model repository [^10] https://pytorch.org/vision/stable/models.html. Nonetheless, in an attempt to facilitate deployment and preserve framework uniformity, these models are converted to TensorFlow, maintaining the same structure and accuracy.
Early Exit (EE) models. For CIFAR-10 models (ResNet-8, ResNet-56, AlexNet), we design early exit models by implementing the methodology presented in (Teerapittayanon et al., 2016). Our approach differs by placing the threshold on the model confidence, i.e., the highest softmax value, instead of the entropy of all the softmax values.
Early exit model implementation follows a two step process. Firstly, a model with the desired architecture is created by adding branches to the original one, which then becomes the backbone of the new model. It possesses one input and as many outputs as branches selected during the design phase (backbone exit also needs to be considered). This model is trained with a joint loss function obtained via a weighted sum over the individual loss functions of each exit. Once the model is successfully trained, the second step involves adding conditional statements for each branch, such that the new model only has one output.
3.3. Performance Metrics and QoS
@@
Minimum accuracy constraint: The accuracy QoS requirement for MNIST and CIFAR-10 image classification is to achieve at $90$%, and for ImageNet-1K is to achieve at least $80$%. Motivated by (Wu et al., 2023), we choose the accuracy requirements $10$% lower than the statte-of-the-art (SOTA) model accuracy for the respective datasets, $99.8$% for MNIST (Mazzia et al., 2021), $99.5$% for CIFAR-10 (Dosovitskiy et al., 2020), and $90$% for ImageNet-1K (Dehghani et al., 2023)).
Maximum latency and energy constraints: The latency QoS requirement is to achieve less than half the time required to offload an image and receive the inference back at the device. The energy QoS requirement is to consume less than half the energy required to offload an image and receive the inference result back at the device.
Note that if latency or energy constraints are relaxed enough to permit offloading an image, remote inference is the best strategy as it achieves higher accuracy. When comparing different strategies in Section 4.4, we fix one of the QoS requirements and determine the strategy that provides the maximum gains for the other two metrics.
3.4. Methodology for Measurements
3.4.1. On-device
All the models examined in this work are either pre-trained by the authors or obtained from existing libraries. Given that INT8 quantized models have lower latency and energy consumption (Hubara et al., 2018), we build and use quantized models albeit at the expense of slightly lower accuracy than the full-precision models. For each device, we load a pre-trained model and measure its latency and energy consumption by performing inference on $n$ images. The averages are then computed, with $n$ varying according to the size of the dataset. Note that, for the inference evaluation, we measure the latency and energy consumption of the API call that initiates the inference routine in each framework.
@@ 3.4.1. On-device
3.4.2. HI
@@
For the HI system, we choose ResNet-8 and ResNet-18 as the local DL models for CIFAR-10 and ImageNet-1K datasets, as both these models are state-of-the-art tinyML models with a reasonable accuracy-to-size ratio. We also train binary LR models (cf. Section 2) for both datasets. Once the inference is performed on the local DL model for each image, its soft-max values are input to the binary LR. The inference is accepted if the binary LR outputs $1$; otherwise, the image is offloaded for remote inference. If the image is set to be offloaded, this is transmitted via the local WiFi network to the edge server. Once the image is fully received on the server, it is preprocessed to the required format of the server model. The server runs the prediction and sends the predicted class back to the device via the network. Given this workflow, the HI system’s accuracy is obtained using (1), where for $i$-th image, $LR_{i}$ denotes the result of the binary LR, $dev\_inf_{i}$ and $ser\_inf_{i}$ denote the device and server predictions, $gt_{i}$ denotes the ground truth class, and $N$ is the size of the test/validation dataset.
(1) $\small Acc_{HI}\!\!=\!\!\frac{\sum_{i=1}^{N}\!\!\mathbbm{1}\!(LR_{i}\!=\!1)\mathbbm{1}\!(dev\_inf_{i}\!=\!gt_{i})\!+\!\mathbbm{1}\!(LR_{i}\!=\!0)\mathbbm{1}\!(ser\_inf_{i}\!=\!gt_{i})}{N}$
Alternatively the accuracy of the HI system can also be represented as
(2) $Acc_{HI}=(Acc_{OD}-\eta_{FN})+(Acc_{S}*\eta_{off})$
where the accuracy obtained by the on-device and remote inference are denoted as $Acc_{OD}$ and $Acc_{S}$ respectively. $\eta_{FN}$ and $\eta_{off}$ represent the fraction of false negative (FN) samples from the binary LR and the fraction of offloaded samples. It can be inferred fron Eq. 2, for a given accuracy QoS, it is important to note that the on-device accuracy does not need to exceed the QoS. The overall system accuracy $Acc_{HI}$ tends to increase with a higher $\eta_{off}$, though this improvement comes at the cost of increased latency and energy consumption. Conversely, the $Acc_{HI}$ decreases as $\eta_{FN}$ increases.
@@
A modular approach was followed for latency and energy measurements for HI. The latency per inference for any image includes the time taken for local DL inference and binary LR decision. If the decision is to offload the image, then the latency accounts for the additional term $t_{off}$, which is the total time taken for image transmission to the server, remote ML inference, and sending back the prediction to the device. Thus, the average latency under HI is given in 3, where $\eta_{off}$ is the fraction of offloaded samples. Similarly, the average on-device energy consumption per image is given in (4), where $e_{edge\_off}$ is the energy consumed to obtain the remote inference for an offloaded image.
(3) $Time_{HI}=t_{dev\_inf}+t_{LR}+\eta_{off}\left[t_{off}\right]$
(4) $Energy_{HI}=e_{dev\_inf}+e_{LR}+\eta_{off}\left[e_{edge\_off}\right]$
3.4.3. EE-HI
@@
For HI system with early exit local DL models, the accuracy, latency, and energy can be computed using (1), (3), and (4). However, in this case, the average latency $t_{device\_inf}$ and the energy consumption $e_{dev\_inf}$ of the early exit DL model depend on the chosen threshold(s) $\theta$ at the early exit branch(s). If the inference of an image satisfies the threshold condition at an early exit branch, then $t_{dev\_inf}$ and $e_{dev\_inf}$ will only include the processing till that branch.
4. Measurement Results and Performance Comparison
4.1. Server Run times
@@
We measure the run times of the remote ML inference on two distinct servers with different GPUs: NVIDIA Tesla T4 and NVIDIA A100. For CIFAR-10, we consider a pre-trained vision transformer model ViT-H/14 (Dosovitskiy et al., 2020), which we fine-tune up to $98.77\%$ accuracy. For ImageNet-1K, we consider a ConvNexT (Liu et al., 2022) with $86.58\%$ accuracy. For the former, we measure an average time per inference of $19.18$ ms for ViT-H/14 (CIFAR-10) and $12.01$ ms for ConvNexT (ImageNet-1K). For the latter, the measured times are $4.29$ ms for ViT-H/14 (CIFAR-10) and $4.41$ ms for ConvNext (ImageNet-1K). These numbers are obtained using half-precision (FP16), a format that uses 16 bits to represent floating-point numbers. Both GPUs have specialized hardware for FP16 arithmetic, which increases performance with negligible accuracy loss. Upon arrival, the image needs to undergo a series of manipulations before it can be passed to the model. Including the entire process, the recorded average time for completion of an inference request becomes 6.95 ms for CIFAR-10 and 5.08 ms for ImageNet-1K. The reason for this is that ConvNeXT uses the same input as the edge models. Therefore, when the image is received by the server, we can immediately perform preprocessing, which takes an average of 0.7 ms. However, preparing data for ViT-H/14 is more involved as it entails the transposition of the input buffer to NCHW format and upscaling to the appropriate dimensions before we can proceed with preprocessing. The additional operations are accountable for the uneven increase in the total execution time compared to ConvNeXT.
@@
Considering that both devices are equipped with NVIDIA hardware, TensorRT is deemed the optimal choice for conducting inference testing. Nonetheless, the size of ViT-H/14 exceeds the 2 GB restriction in the Protobuf (Protocol Buffers) serialized message size  [^14] https://protobuf.dev/programming-guides/encoding/, rendering it impossible to parse the model into the framework. Instead, the ONNX Runtime was utilized with the CUDA execution provider, given that TensorRT execution provider naturally fails in this framework for the same reason. However, the ConvNeXT model did not have this issue, and the inference was executed using TensorRT execution provider within ONNX Runtime.
@@
Table 3. Latency and energy consumption to offload an image using WiFi.
@@
Table 4. Accuracy, latency and energy results for Coral Micro, Raspberry Pi and Jetson Orin devices
@@
4.2. Wireless Communication Timings
4.2.1. Image transmission: WiFi vs BLE
Given the availability of both BLE and Wi-Fi for all the devices (except Arduino Nano), the question of which one should be used arises. To address this, an experiment was performed on a Raspberry Pi, where both latency and energy were measured for transmitting a CIFAR-10 image to the server and waiting for a response. It is worth noting that the energy measurements were taken for the whole working device, not only the communication module.

### v2

@@ III-A Devices
In this study, we conduct measurements on five devices shown in Fig. 3 and described in Tables I and II.
@@
Arduino Nano 33 BLE Sense and ESP32 (cf. Table I) are the least powerful devices used in this paper. On-device inference is performed using the TensorFlow Lite Micro framework, which requires loading the model as a binary file, leading to slower on-device computation. Furthermore, in contrast to all other devices, Arduino Nano does not support Wi-fi communication. For this reason, offloaded images are transmitted to the edge server via Bluetooth Low Energy (BLE).
@@
Coral Dev Board Micro (cf. Table I) posseses higher processing power and more memory than other commodity MCUs. It can load and execute large compute-intensive DL models with higher inference accuracy. Although the device has a built-in TPU, in certain cases, it falls short. For example, the TPU supports a limited set of instructions [^4] https://coral.ai/docs/edgetpu/models-intro/#supported-operations, which directly impact the models it can execute using the TPU. The TFLite models must be compiled using edgetpu-compiler to be executed on the TPU. Further, because EE models use an intermediate layer with an if instruction, they cannot be compiled to a model supported by the TPU because the instruction is not supported. One potential mitigation of such an issue is running the model on the CPU at the expense of increased inference latency and energy consumption.
@@
Contrary to Jetson Orin or Coral Micro, the Raspberry Pi 4B (cf. Table II) does not benefit from any hardware-based AI acceleration. Such selection provides us with a broader view of the available system-on-chip ecosystem. Regarding the software implementation, all the models are loaded and executed using the TFLite Python API. In this framework, all the necessary operations for the tested models are supported [^5] https://www.tensorflow.org/mlir/tfl_ops. Combined with the fact that the models were run on the CPU, this enables smooth deployment, even for Early Exit implementations.
@@
Jetson Orin Nano (cf. Table II) includes an integrated GPU that features Tensor Cores. For this device, we have used TensorRT SDK [^6] https://docs.nvidia.com/deeplearning/tensorrt, which is a framework developed by NVIDIA for high-performance deep learning inference, built on top of CUDA. It optimizes the models using layer and tensor fusion and kernel tuning to maximize performance on the specific hardware of the target device. To port the pre-trained models to TensorRT, we must first convert them to ONNX. This pipeline yields correct results for base models, but for early exit we find that the TensorRT engine fails to produce a graph-capturable network in CUDA, which causes early exit models to be slower than the base implementations. To mitigate this issue, transforming the Tensorflow subgraph models to PyTorch successfully generates models that effectively utilize CUDA graphs and achieve the required performance. Jetson Orin has two power modes: one for maximum performance (15W) with all hardware and clocks at full capacity and another for low power consumption (7W) with RAM, CPU, GPU, and other components capped. We examine both power modes to investigate their effects and observe that enabling dynamic clock rate results in higher latency and variability for relatively small energy savings. Setting the clock to maximum in each power mode results in higher performance per watt and more stable outcomes under constant load.
Figure 3: Devices used in the study. Starting from the left: Arduino Nano 33 BLE Sense, ESP32 DevKit-C, Coral Dev Board Micro, Raspberry Pi 4B, Jetson Orin Nano.
III-B Datasets, Models and Training
We study image classification tasks using three datasets MNIST, CIFAR-10, and ImageNet-1K. As discussed in Section I, MNIST image classification tasks can be performed on Aruduino Nano and ESP32, without offloading to a remote ML. Thus, we focus on CIFAR-10 and ImageNet-1K image classification tasks in the rest of the paper. The details of the datasets are presented in Table III.
TABLE I: Device characteristics of MCUs used for the measurements
@@ III-B Datasets, Models and Training
TABLE II: Device characteristics of the moderately powerful devices used for the measurements
@@
We use TensorFlow [^7] https://www.tensorflow.org/ (specifically, TFLite [^8] https://www.tensorflow.org/lite and TFLite Micro [^9] https://www.tensorflow.org/lite/microcontrollers) as the DL software library for most of this work due to its popularity and convenience in compiling models for on-device deployment. We refer to the models we introduce below as base models to differentiate them from the models we design and train using early exit branches, later in Section VII.
@@
ImageNet-1K. For ImageNet, we select ResNet-18, ResNet-50, AlexNet, and RegNetY32GF [60]. All these models are retrieved from PyTorch’s own model repository [^11] https://pytorch.org/vision/stable/models.html. Nonetheless, in an attempt to facilitate deployment and preserve framework uniformity, these models are converted to TensorFlow, maintaining the same structure and accuracy.
Early Exit (EE) models. We design on-device early exit models for CIFAR-10 models (ResNet-8, ResNet-56, AlexNet) by implementing the methodology presented in [24]. Our approach differs from [24] in that we use a threshold on the model confidence, i.e., the maximum softmax value, instead of the entropy of all the softmax values. Later in Section VII, we present the details of our design of the early exit models.
TABLE III: Details of the datasets used
Dataset Number of Samples Input Size Number of Classes
MNIST 70,000 (60,000 train, 10,000 test) $28\times 28$ grayscale 10
CIFAR-10 60,000 (50,000 train, 10,000 test) $32\times 32\times 3$ RGB 10
ImageNet-1K $\sim$1.28M (train), 50,000 (val), 100,000 (test) $224\times 224\times 3$ RGB (varies) 1,000
III-C Performance Metrics and QoS
@@
Accuracy QoS constraint: Motivated by [61], we choose the accuracy requirements $10$% lower than the state-of-the-art (SOTA) model accuracy for the respective datasets, $99.8$% for MNIST [62], $99.5$% for CIFAR-10 [63], and $90$% for ImageNet-1K [64]). Thus, the accuracy QoS requirement for MNIST and CIFAR-10 image classification is to achieve at $90$%, and for ImageNet-1K is to achieve at least $80$%.
Latency and energy constraints: The average latency QoS requirement in this context is defined dynamically, requiring inference to be completed in less than half the time needed to offload an image and receive the result back at the device. Similarly, the average energy QoS requirement is dynamically set to use less than half the energy required for the same offloading process.
Note that if latency or energy constraints are relaxed enough to permit offloading an image, remote inference is the best strategy as it achieves higher accuracy. When comparing different strategies in Section VI, we fix one of the QoS requirements and determine the strategy that provides the maximum gains for the other two metrics.
IV Methodology for Measurements
IV-1 On-device
All the models examined in this work are either pre-trained by the authors or obtained from existing libraries. Given that INT8 quantized models have lower latency and energy consumption [65], we build and use quantized models albeit at the expense of slightly lower accuracy than the full-precision models. For each device, we load a pre-trained model and measure its latency and energy consumption by performing inference on $n$ images, where $n$ is the size of the validation dataset. The averages are then computed, with $n$ varying according to the size of the validation dataset. Note that, for the inference evaluation, we measure the latency and energy consumption of the API call that initiates the inference routine in each framework.
@@ IV-1 On-device
IV-2 HI
For the HI system, we choose ResNet-8 and ResNet-18 as the on-device DL models for CIFAR-10 and ImageNet-1K datasets, as both are tinyML models with a reasonable accuracy-to-size ratio.
The binary LR algorithm: For the HI decision-making, we use a binary LR algorithm for each dataset. We train the LR using the soft-max values output by the on-device model for each image (from the training set) as the input. The output labels during the training are ${0,1}$, where 00 corresponds to on-device inference being incorrect and $1$ corresponds to the inference being correct.
@@
Specifically, in the testing phase, for the $i$-th image, the inference is accepted if the binary LR outputs $1$, denoted by $LR_{i}=1$; otherwise, the image is offloaded for remote inference, denoted by $LR_{i}=0$. If the image is set to be offloaded, this is transmitted via the local WiFi network (or Bluetooth in case of Arduino Nano), to the edge server. Once the image is received on the server, it is preprocessed to the format required at the server model. The server processes the prediction and sends the predicted class back to the device via the network. To quantify the accuracy of the HI system, we define key terms: let $dev\_inf_{i}$ and $ser\_inf_{i}$ denote the device and server predictions, and $gt_{i}$ denotes the ground truth class for image $i$. The HI system’s accuracy, denoted by $Acc_{HI}$, is given by,
$Acc_{HI}\!\!=\!\!\frac{1}{n}\sum_{i=1}^{n}\!\Big{[}\!\mathbbm{1}\!(LR_{i}\!=\!1)\mathbbm{1}\!(dev\_inf_{i}\!=\!gt_{i})$
$\quad\quad\quad\quad\quad\quad+\mathbbm{1}\!(LR_{i}\!=\!0)\mathbbm{1}\!(ser\_inf_{i}\!=\!gt_{i})\Big{]}$ (1)
$=\!\!\frac{1}{n}\sum_{i=1}^{n}\!\Big{[}\!\mathbbm{1}\!(dev\_inf_{i}\!=\!gt_{i})\!-\!\mathbbm{1}\!(LR_{i}\!=\!0)\mathbbm{1}\!(dev\_inf_{i}\!=\!gt_{i})$
$\quad\quad\quad\quad\quad\quad+\mathbbm{1}\!(LR_{i}\!=\!0)\mathbbm{1}\!(ser\_inf_{i}\!=\!gt_{i})\Big{]}$
$=\!Acc_{OD}\!-\!\eta_{FN}\!+\!\frac{1}{n}\!\sum_{i=1}^{n}\!\!\mathbbm{1}\!(LR_{i}\!=\!0)\mathbbm{1}\!(ser\_inf_{i}\!=\!gt_{i}),$ (2)
where $\mathbbm{1}(\cdot)$ is an indicator function, $Acc_{OD}$ represents the accuracy of the on-device model, and $\eta_{FN}$ denotes the fraction of false negatives under the binary LR, i.e., the fraction of data samples offloaded by the binary LR that are correctly classified by the on-device model. From (2), we infer that while the on-device model’s accuracy may fall below the accuracy QoS requirement, $Acc_{HI}$ can still meet the requirement if the remote model’s accuracy is significantly higher than the QoS requirement and $\eta_{FN}$ is sufficiently small (ideally zero). This underscores the critical role of the binary LR’s performance in HI decision-making. From the above observations, we conclude that employing an HI system with a smaller on-device model, even with lower accuracy, can lead to improved overall performance. It is worth noting, however, that a naive HI decision-making algorithm could reduce $\eta_{FN}$ by increasing the number of offloads, thereby boosting overall system accuracy $Acc_{HI}$. Yet, this approach comes at the cost of increased false positives, as well as higher latency and energy consumption per inference.
@@
We follow a modular approach for latency and energy measurements for HI. The latency per inference for an image includes the processing time for the on-device model and the binary LR decision, denoted by $t_{dev\_inf}$ and $t_{LR}$, respectively. Let $e_{dev\_inf}$ and $e_{LR}$ denote the respective energy consumptions. If the decision is to offload the image, then the latency accounts for the additional term $t_{off}$, which is the total time taken for image transmission to the server, remote ML inference, and sending back the prediction to the device. Similarly, the additional device energy consumption for an offloaded image is denoted by $e_{off}$, which includes the transmission energy. Let $\eta_{off}$ denote the fraction of offloaded samples in the HI system, then the average time per inference, $Time_{HI}$, and average energy per inference, $Energy_{HI}$, are given by
$Time_{HI}$ $=t_{dev\_inf}+t_{LR}+\eta_{off}\left[t_{off}\right]$ (3)
$Energy_{HI}$ $=e_{dev\_inf}+e_{LR}+\eta_{off}\left[e_{off}\right].$ (4)
IV-3 EE-HI
@@
For HI system with early exit on-device models, the accuracy, latency, and energy can be computed using (1), (3), and (4). However, in this case, the average latency $t_{dev\_inf}$ and the energy consumption $e_{dev\_inf}$ of the early exit DL model depend on the chosen threshold(s) $\theta$ at the early exit branch(s). If the inference of an image satisfies the threshold condition at an early exit branch, then $t_{dev\_inf}$ and $e_{dev\_inf}$ will only include the processing till that branch.
V Measurement Results
V-A Server Run times
@@
We measure the run times of the remote ML inference on two distinct servers with different GPUs: NVIDIA Tesla T4151515https://www.nvidia.com/en-us/data-center/tesla-t4/ and NVIDIA A100161616https://www.nvidia.com/en-us/data-center/a100/. The comparison of the computing powers of these servers are presented in Table IV. For CIFAR-10, we consider a pre-trained vision transformer model ViT-H/14 [63], which we fine-tune to achieve an accuracy of $98.77\%$. For ImageNet-1K, we consider a ConvNexT [66] with $86.58\%$ accuracy. For NVIDIA Tesla T4, we measure an average time per inference of $19.18$ ms for ViT-H/14 (CIFAR-10) and $12.01$ ms for ConvNexT (ImageNet-1K). For NVIDIA A100, the measured times are $4.29$ ms for ViT-H/14 (CIFAR-10) and $4.41$ ms for ConvNext (ImageNet-1K). These numbers are obtained using half-precision (FP16), a format that uses 16 bits to represent floating-point numbers. Both GPUs have specialized hardware for FP16 arithmetic, which increases performance with negligible accuracy loss. Upon arrival, the image needs to undergo pre-processing, including conversion of image dimensions to match the input dimensions of the model. We measure the end-to-end average processing time of an inference request as 6.95 ms for CIFAR-10 and 5.08 ms for ImageNet-1K. The reason for smaller inference time for ImageNet-1k images is that ConvNeXT uses the same dimension image as input as the edge models and does not require converting the image dimensions. Therefore, when the image is received by the server, we can immediately perform preprocessing, which takes an average of 0.7 ms. However, preparing data for ViT-H/14 is more involved, as it requires transposing the input buffer to the NCHW format [^17] https://docs.nvidia.com/deeplearning/cudnn/latest/developer/core-concepts.html and upscaling to the appropriate dimensions before preprocessing. The additional operations are accountable for the uneven increase in the total execution time compared to ConvNeXT.
@@
Considering that both servers are equipped with NVIDIA hardware, TensorRT is deemed the optimal choice for conducting inference testing. Nonetheless, the size of ViT-H/14 exceeds the 2 GB restriction in the Protobuf (Protocol Buffers) serialized message size [^18] https://protobuf.dev/programming-guides/encoding/, rendering it impossible to parse the model into the framework. Instead, the ONNX Runtime was utilized with the CUDA execution provider, given that TensorRT execution provider naturally fails in this framework for the same reason. However, the ConvNeXT model did not have this issue, and the inference was executed using TensorRT execution provider within ONNX Runtime.
@@
TABLE IV: Computing Resource Comparison of Edge Servers
Feature NVIDIA Tesla T4 NVIDIA A100
Architecture Turing Ampere
Memory 16GB GDDR6 80GB HBM2e
Memory Bandwidth 320 GB/s 1,555 GB/s
CUDA Cores 2,560 6,912
Tensor Cores 320 432
TABLE V: Latency (communication + server time) and energy consumption to offload an image using WiFi.
@@
TABLE VI: Accuracy, latency and energy results for Coral Micro, Raspberry Pi and Jetson Orin devices
@@
V-B Wireless Communication Timings
V-B1 Image transmission – WiFi vs BLE